OpenStudy (anonymous):

Integrate: xcotx^(2) dx

OpenStudy (anonymous):

\[xcotx^2 dx\]

hartnn (hartnn):

did you try the substitution x^2 = y ?

OpenStudy (anonymous):

like u-substitution?

hartnn (hartnn):

yeah, so, x^2 =u what's du=.... ?

OpenStudy (anonymous):

2x

hartnn (hartnn):

more precisely its du = 2x dx right ? so, x dx = (du)/2 put this in your integral, what you get as new integral ?

OpenStudy (anonymous):

(1/2u)(cotu)

OpenStudy (anonymous):

oh okay so the final answer would be 1/2ln(sinx^(2))+c?

hartnn (hartnn):

oh, did you mean (1/2)*du*cot u there ? and yes, thats correct.

OpenStudy (anonymous):

yeah sorry. i meant that.

hartnn (hartnn):

1/2ln(sinx^(2))+c is correct :)

OpenStudy (anonymous):

thanks for your help on this!

hartnn (hartnn):

you are most welcome ^_^